OpenStudy (anonymous):

find the perimeter of the rectangle 3x+6 5x-2

OpenStudy (shamil98):

I'm guessing those are the values of the length of width. If you recall the perimeter of a rectangle is \[\large P = 2L + 2W\]

OpenStudy (anonymous):

Shamil98 that is the right equation.

OpenStudy (anonymous):

Thanks for commenting. take a look at the attached file

OpenStudy (shamil98):

Ah, okay.

OpenStudy (shamil98):

First, off, we want to solve for x.

OpenStudy (anonymous):

Thank you.

OpenStudy (shamil98):

\[\large 3x + 6 = 5x -2\] \[\large 8 = 2x\] \[\large 4 = x\]

OpenStudy (shamil98):

Evaluate the width. 3(4) + 6 12 + 6 w = 18 Now back to your equation for perimeter. p = 2l + 2w p = 2(6) + 2(18) simplify to get your final result.

OpenStudy (anonymous):

THank you shami... can you take a look at this attachment as well..

OpenStudy (anonymous):

what does the last part say? 4x+?

OpenStudy (anonymous):

4x+9

OpenStudy (shamil98):

All sides are equal. so. 9x - 11= 4x + 9 solve for x, and the evaluate either expression.

OpenStudy (anonymous):

i'm getting that wrong... please help me solve for x

OpenStudy (shamil98):

9x - 11= 4x + 9 5x = 20 x = 4 4(4) + 9 = 16 + 9 = 25

OpenStudy (shamil98):

The perimeter is all the sides added. triangles have three sides. all sides are equal. 25 x 3 = 75

OpenStudy (anonymous):